# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance (Extended version)

### v3

@@ I INTRODUCTION
Safe autonomous robot navigation consists in steering a robot to a target location while avoiding obstacles. One common method is the Navigation Function (NF) approach [1], where a combination of attractive and repulsive vector fields, with a properly tuned parameter, guarantees safe, almost [^1] Almost global convergence refers to convergence from all initial locations except a set of zero Lebesgue measure. global convergence of the robot to the target location in sphere worlds. For general convex and star-shaped obstacles, diffeomorphic mappings [1, 2] allow the use of the NF approach but require complete knowledge of the environment. In [3], the NF method has been extended to convex obstacles that satisfy certain curvature properties, and are not too flat and not too close to the target. The work in [4] removes this flatness limitation, in environments with ellipsoidal obstacles, by using the Hessian to locally transform areas near obstacles into spherical regions. However, similar to [3], it is assumed that the entire shape of the obstacle is known when the robot visits its neighbourhood.
@@
In [10], a feedback controller based on Nagumo’s theorem [11, Theorem 4.7] was proposed for autonomous navigation around general obstacles. The method ensures forward invariance of the obstacle-free space by projecting the ideal velocity control vector, which points towards the target, onto the tangent cone at the obstacle boundary when necessary. The research in [10] was extended in [12] to guarantee almost global asymptotic stability in unknown environments with strongly convex obstacles. The work in [13] introduced a reactive navigation approach using separating hyperplanes for robots in environments with unknown, sufficiently separated, strongly convex obstacles. This was further extended in [14] to handle partially known environments with non-convex obstacles, where the robot has information about the shape of the non-convex obstacles but not their locations in the workspace.
However, the approaches discussed above provide, at best, almost global convergence guarantees due the undesired equilibria that are generated when using continuous time-invariant vector fields [15]. This can be resolved by introducing discontinuities in the control, as shown in [16, 17, 18, 19, 20].
In [16], a discontinuous feedback law was developed for robot navigation in partially known two-dimensional environments, guiding the robot with the navigation function gradient near known obstacles and boundary-following near unknown ones. In [17], the authors proposed a discontinuous feedback controller for autonomous navigation of nonholonomic robots in two-dimensional environments with non-convex obstacles, imposing restrictions on the curvatures of obstacles’ boundaries and the inter-obstacle arrangements. In [18], hybrid control methods are employed to achieve robust global asymptotic convergence of a group of planar robots to the maximizer of an unknown potential field, whose value can be measured at any location, while avoiding a single obstacle in a 2-dimensional environment. In [19], the authors proposed a hybrid control law to globally asymptotically stabilize a class of linear systems with drift while avoiding neighbourhoods of unsafe isolated points. In [20], hybrid control techniques were employed to achieve global stabilization of target locations in $n$-dimensional environments with sufficiently separated ellipsoidal obstacles.
@@ II-A Notations
For two sets $\mathcal{A},\mathcal{B}\subset\mathbb{R}^{n}$, the relative complement of $\mathcal{B}$ with respect to $\mathcal{A}$ is denoted by $\mathcal{A}\setminus\mathcal{B}=\{\mathbf{a}\in\mathcal{A}|\mathbf{a}\notin\mathcal{B}\}$. The symbols $\partial\mathcal{A},\mathcal{A}^{\circ}$, $\mathcal{A}^{c}$ and $\bar{\mathcal{A}}$ represent the boundary, interior, complement and the closure of the set $\mathcal{A}$, respectively, where $\partial\mathcal{A}=\bar{\mathcal{A}}\backslash\mathcal{A}^{\circ}$. The cardinality of a set $\mathcal{A}$ is denoted by $\mathbf{card}(\mathcal{A})$. The Minkowski sum of the sets $\mathcal{A}$ and $\mathcal{B}$ is denoted by, $\mathcal{A}\oplus\mathcal{B}=\{\mathbf{a}+\mathbf{b}|\mathbf{a}\in\mathcal{A},\mathbf{b}\in\mathcal{B}\}$. The dilated version of a set $\mathcal{A}\subset\mathbb{R}^{n}$ with $r\geq 0$ is represented by $\mathcal{D}_{r}(\mathcal{A})=\mathcal{A}\oplus\mathcal{B}_{r}(\mathbf{0})$. The $r-$neighbourhood of a set $\mathcal{A}$ is denoted by $\mathcal{N}_{r}(\mathcal{A})=\mathcal{D}_{r}(\mathcal{A})\setminus\mathcal{A}^{\circ}$ where $r$ is a strictly positive scalar.
@@ II-D Hybrid system framework
[…]
A subset $\mathbb{T}\subset\mathbb{R}_{\geq}\times\mathbb{N}$ is a hybrid time domain if it is a union of a finite or infinite sequence of intervals $[t_{j},t_{j+1}]\times\{j\},$ where the last interval (if existent) is possibly of the form $[t_{j},T)$ with $T$ finite or $T=+\infty$. The ordering of points on each hybrid time domain is such that $(t,j)\preceq(t^{\prime},j^{\prime})$ if $t\leq t^{\prime}$ and $j\leq j^{\prime}$. A hybrid solution $\phi$ is maximal if it cannot be extended, and complete if its domain dom $\phi$ (which is a hybrid time domain) is unbounded.
@@ III PROBLEM FORMULATION
We consider a spherical robot with radius $r\geq 0$ operating in a convex, three-dimensional Euclidean space $\mathcal{W}\subseteq\mathbb{R}^{3}.$ The workspace $\mathcal{W}$ is cluttered with finite number of compact, convex obstacles $\mathcal{O}_{i}\subset\mathcal{W},i\in\{1,\ldots,b\},b\in\mathbb{N}.$ We define obstacle $\mathcal{O}_{0}:=(\mathcal{W}^{\circ})^{c}$ as the complement of the interior of the workspace. Collectively, the obstacle-occupied workspace is denoted by $\mathcal{O}_{\mathcal{W}}=\bigcup_{i\in\mathbb{I}}\mathcal{O}_{i}$, where $\mathbb{I}=\{0,\ldots,b\}.$ We make the following workspace feasibility assumption:
@@ IV-B Geometric construction of the flow and jump sets
When the robot operates in the move-to-target mode, its velocity is directed towards the target location. Hence, if the path joining the robot’s location and the target is obstructed by an obstacle, let’s say $\mathcal{O}_{i}$ for some $i\in\mathbb{I}$ i.e., $\mathcal{L}_{s}(\mathbf{x},\mathbf{0})\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\neq\emptyset$, then the center of the robot enters the $(r_{a}+\gamma)-$neighbourhood of the obstacle $\mathcal{O}_{i}$ through the landing region $\mathcal{R}_{l}^{i},$ which is defined as
@@ IV-B1 Flow and jump sets (move-to-target mode)
where $\gamma_{s}\in(0,\gamma).$ For robustness purposes (with respect to noise), we introduce a hysteresis region by allowing the robot, operating in the move-to-target mode inside the $(r_{a}+\gamma)-$neighbourhood of the set $\mathcal{O}_{\mathcal{W}}$, to move closer to the set $\mathcal{O}_{\mathcal{W}}$ before switching to the obstacle-avoidance mode.
@@ IV-B2 Flow and jump sets (obstacle-avoidance mode)
The robot operates in the obstacle-avoidance mode inside the $\gamma-$neighbourhood of the obstacle-occupied workspace. Since the robot can safely move straight towards the target location and exit the $\gamma-$neighborhood of the nearest obstacle, whenever it is in the exit region (17), it should switch back to the move-to-target mode only if it is in the exit region. To that end, we make use of the hit point $\mathbf{h}$ (i.e., the location of the center of the robot when it switched from the move-to-target mode to the current obstacle-avoidance mode) to define the jump set of the obstacle-avoidance mode $\mathcal{J}_{1}^{\mathcal{W}}$ for the state $\mathbf{x}$ as follows:
@@ IV-C Update law $\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m,s)$
Control design summary: The proposed hybrid feedback control law can be summarized as follows:
  • Parameters selection: the target location is set at the origin with $\mathbf{0}\in\mathcal{W}_{r_{a}}^{\circ}$, and $\xi(0,0)\in\mathcal{K}$. The gain parameters $\kappa_{s}$ and $\kappa_{r}$ are set to positive values, and sufficiently small positive value is chosen for $\bar{\epsilon}$, used in (22). The scalar parameter $\gamma$, used in the construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, is selected such that $\gamma\in(0,\delta-r_{s})$. The parameters $\gamma_{a}$ and $\gamma_{s}$ are set to satisfy $0<\gamma_{a}<\gamma_{s}<\gamma.$ The parameter $\delta_{s}$, used in (24), is chosen such that $\delta_{s}\in(0,\tau_{s})$, where $\tau_{s}>0$.
  • Move-to-target mode $m=0$: this mode is activated when $\xi\in\mathcal{F}_{0}$. As per (11a), the control input is given by $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x}$, causing $\mathbf{x}$ to evolve along the line segment $\mathcal{L}_{s}(\mathbf{0},\mathbf{x})$ towards the origin. If, at some instance of time, $\xi$ enters in the jump set $\mathcal{J}_{0}$ of the move-to-target mode, the state variables $(\mathbf{h},\mathbf{a},m,s)$ are updated using (27), and the control input switches to the obstacle-avoidance mode.
  • Obstacle-avoidance mode $m=1$: this mode is activated when $\xi\in\mathcal{F}_{1}$. As per (11a), the control input is given by $\mathbf{u}(\xi)=\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{a})$, causing $\mathbf{x}$ to evolve in the $\gamma-$neighborhood of the nearest obstacle along the hyperplane $\mathcal{P}(\mathbf{h},\mathbf{a})$ until the state $\xi$ enters in the jump set $\mathcal{J}_{1}$ of the obstacle-avoidance mode. When $\xi\in\mathcal{J}_{1}$, the state variables $(m,s)$ are updated using (30), and the control input switches to the move-to-target mode.
@@ V STABILITY ANALYSIS
The next lemma shows that the hybrid closed-loop system (31) satisfies the hybrid basic conditions [23, Assumption 6.5], which guarantees the well-posedness of the hybrid closed-loop system.
@@ VII SIMULATION RESULTS
In the first simulation, we consider a single three-dimensional convex obstacle, as shown in Fig. 2. The initial location of the robot $\mathbf{x}(0,0)$ is represented by the green diamond symbol, and the origin is selected as the target location. The radius of the robot $r=0.15m$ and the safety distance $r_{s}=0.1m$. The parameter $\gamma=0.4m$ and the parameter $\epsilon$, used in (22), is set to $0.1m$. The gains $\kappa_{s}$ and $\kappa_{r}$, used in (11a), are set to $1$ and $0.5$, respectively. The sensing radius $R_{s}$, used in (33), is set to $2m$. The simulations are performed in MATLAB 2020a. In Fig. 2, the red-colored trajectory represents the motion of the robot in the move-to-target mode, whereas the blue-colored trajectory represents the motion of the robot in the obstacle-avoidance mode. When the robot switches from the move-to-target mode to the obstacle-avoidance mode, the algorithm updates the coordinates of the hit point $\mathbf{h}$ and the unit vector $\mathbf{a}$ according to (27). Then, the robot moves in the $\gamma-$neighborhood of the nearest obstacle with its center in the hyperplane $\mathcal{P}(\mathbf{h},\mathbf{a})$, as shown in Fig. 2. When the center of the robot enters in the jump set of the obstacle-avoidance mode, it switches back to the move-to-target mode and converges to the target location at the origin.
Figure 2: The trajectory of the robot safely converging to the target location at the origin.
We consider an unbounded workspace i.e., obstacle $\mathcal{O}_{0}=\emptyset$, with $2$ three-dimensional, convex obstacles, as shown in Fig. 3. We apply the proposed hybrid feedback controller (11) for robot initialized at $8$ different locations in the obstacle-free workspace. The target is located at the origin. The radius of the robot is set to $0.15m$. The minimum safety distance $r_{s}=0.05m$ and the parameter $\gamma=0.2m$. The gains $\kappa_{s}$ and $\kappa_{r}$, used in (11a), are set to $0.5$ and $1$, respectively. The sensing radius $R_{s}$, used in (33), is set to $1m$. The parameter $\epsilon$, used in (22), is set to 0.05m. In Fig. 3, the dotted portion of the trajectory represents the motion of the robot in the move-to-target mode, whereas the solid portion corresponds to the movement in the obstacle-avoidance mode. From Fig. 3, it can be observed that the robot converges to the target location while simultaneously avoiding collisions with the obstacles. Fig. 4 shows that the center of the robot stays at least $r_{a}$ meters away from the boundary of the obstacles.
Figure 3: Robot trajectories starting from different locations.
Figure 4: Distance of the center of the robot $\mathbf{x}$ from the boundary of the obstacle occupied workspace $\mathcal{O}_{\mathcal{W}}$.
We compare our approach with the separating hyperplane method developed in [13] which can be implemented in a priori unknown environments using information from a range-bearing sensor mounted on the robot. However, unlike our approach, it only works for strongly convex obstacles [13, Assumption 2]. The workspace contains two obstacles, an ellipsoidal obstacle (strongly convex) and a polygonal obstacle that does not satisfy [13, Assumption 2]. The robot’s radius is set to $0.2m$, with a minimum safety distance $r_{s}=0.1m$ and $\gamma=0.5m$. The gains $\kappa_{s}=0.4$ and $\kappa_{r}=2$ are used in (11a), while the sensing radius $R_{s}=2m$ is used in (33), and $\epsilon=0.5m$ in (22).
@@ IX-B Proof of Lemma 2
IX-C Proof of Lemma 3
First we prove that the union of the flow and jump sets covers exactly the obstacle-free state space $\mathcal{K}$. For $m=0$, according to (18) and (19), by construction we have $\mathcal{F}_{0}^{\mathcal{W}}\cup\mathcal{J}_{0}^{\mathcal{W}}=\mathcal{W}_{r_{a}}.$ Similarly, for $m=1$, according to (21) and (23), by construction one has $\mathcal{F}_{m}^{\mathcal{W}}\cup\mathcal{J}_{m}^{\mathcal{W}}=\mathcal{W}_{r_{a}}$. Inspired by [20, Appendix 11], the satisfaction of the following equation:
$\mathcal{F}_{m}^{\mathcal{W}}\cup\mathcal{J}_{m}^{\mathcal{W}}=\mathcal{W}_{r_{a}},m\in\mathbb{M},$ (35)
along with (20), (24) and (25) implies $\mathcal{F}\cup\mathcal{J}=\mathcal{K}.$
Now, inspired by [20, Appendix 1], for the hybrid closed-loop system (31), with data $\mathcal{H}=(\mathcal{F},\mathbf{F},\mathcal{J},\mathbf{J}),$ define $\mathbf{S}_{\mathcal{H}}(\mathcal{K})$ as the set of all maximal solutions $\xi$ to $\mathcal{H}$ with $\xi(0,0)\in\mathcal{K}$. Since $\mathcal{F}\cup\mathcal{J}=\mathcal{K}$, each $\xi\in\mathbf{S}_{\mathcal{H}}(\mathcal{K})$ has range rge $\text{rge }\xi\subset\mathcal{K}.$ Additionally, if every maximal solution $\xi\in\mathbf{S}_{\mathcal{H}}(\mathcal{K})$ is complete, then the set $\mathcal{K}$ will be forward invariant [25, Definition 3.13]. Since the hybrid closed-loop system (31) satisfies the hybrid basic conditions, as stated in Lemma 2, one can use [23, Proposition 6.10], to verify the following viability condition:
@@ IX-C Proof of Lemma 3
which will allow us to establish the completeness of the solution $\xi$ to the hybrid closed-loop system (31). In (36), $\mathbf{T}_{\mathcal{F}}(\xi)$ represents the tangent cone to the set $\mathcal{F}$ at $\xi$.
Let $(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{F}\setminus\mathcal{J}$, which implies by (20), (24) and (25) that $(\mathbf{x},\mathbf{h},\mathbf{a},s)\in(\mathcal{F}_{m}^{\mathcal{W}}\setminus\mathcal{J}_{m}^{\mathcal{W}})\times\mathcal{W}_{r_{a}}\times\mathbb{S}^{2}\times\mathbb{R}_{\geq 0}$ for some $m\in\mathbb{M}$. For $\mathbf{x}\in(\mathcal{F}_{m}^{\mathcal{W}})^{\circ}\setminus\mathcal{J}_{m}^{\mathcal{W}}$ with $(\mathbf{h},\mathbf{a},m,s)\in\mathcal{W}_{r_{a}}\times\mathbb{S}^{2}\times\mathbb{M}\times\mathbb{R}_{\geq 0}$ the tangent cone $\mathbf{T}_{\mathcal{F}}(\xi)=\mathbb{R}^{3}\times\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})\times\mathcal{P}(\mathbf{0},\mathbf{a})\times\{0\}\times\mathbf{T}_{\mathbb{R}_{\geq 0}}(s),$ where the set $\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$ is given by
$\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})=\begin{cases}\begin{matrix}[l]\mathbb{R}^{3},&\text{if }\mathbf{h}\in(\mathcal{W}_{r_{a}})^{\circ},\\ \mathcal{P}_{\geq}(\mathbf{0},(\mathbf{h}-\Pi(\mathbf{h},\mathcal{O}_{\mathcal{W}}))),&\text{if }\mathbf{h}\in\partial\mathcal{W}_{r_{a}},\end{matrix}\end{cases}$ (37)
where for $\mathbf{h}\in\partial\mathcal{W}_{r_{a}}$, the projection $\Pi(\mathbf{h},\mathcal{O}_{\mathcal{W}})$ is unique. For $s\in\mathbb{R}_{\geq 0}$, the set $\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ is defined as
$\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)=\begin{cases}\begin{matrix}[l]\mathbb{R},&\text{if }s\in\mathbb{R}_{>0},\\ \mathbb{R}_{\geq 0},&\text{if }s=0.\end{matrix}\end{cases}$ (38)
Since, according to (31), $\dot{\mathbf{h}}=\mathbf{0}$ and $\dot{s}=1$, we have $\dot{\mathbf{h}}\in\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$ and $\dot{s}\in\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$, respectively, and (36) holds.
Next, we consider the case where $\xi\in\mathcal{F}\setminus\mathcal{J}$ with $\mathbf{x}\in\partial\mathcal{F}_{m}^{\mathcal{W}}\setminus\mathcal{J}_{m}^{\mathcal{W}},m\in\mathbb{M}$. For $m=0$, according to (18) and (19), one has
$\partial\mathcal{F}_{0}^{\mathcal{W}}\setminus\mathcal{J}_{0}^{\mathcal{W}}\subset\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}})\cap\mathcal{R}_{e},$ (39)
and for $\mathbf{x}\in\partial\mathcal{F}_{0}^{\mathcal{W}}\setminus\mathcal{J}_{0}^{\mathcal{W}},$ the projection $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ is unique. Hence, for all $\xi\in\mathcal{F}_{0}\setminus\mathcal{J}_{0}$ with $\mathbf{x}\in\partial\mathcal{F}_{0}^{\mathcal{W}}\setminus\mathcal{J}_{0}^{\mathcal{W}}$,
$\mathbf{T}_{\mathcal{F}}(\xi)=\mathcal{P}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\times\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})\times\mathcal{P}(\mathbf{0},\mathbf{a})\times\{0\}\times\mathbf{T}_{\mathbb{R}_{\geq 0}}(s),$ (40)
where $\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$ and $\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ are defined in (37) and (38), respectively, and $\mathbf{x}_{\pi}=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$. Also, according to (11a), for $m=0$, one has $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x},\kappa_{s}>0$. According to (17) and (39), for $\xi\in\mathcal{F}_{0}\setminus\mathcal{J}_{0}$ with $\mathbf{x}\in\partial\mathcal{F}_{0}^{\mathcal{W}}\setminus\mathcal{J}_{0}^{\mathcal{W}}$, one can conclude that $\mathbf{u}(\xi)\in\mathcal{P}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})$. Moreover, according to (31), it is clear that $\dot{\mathbf{h}}=\mathbf{0}\in\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$, $\dot{s}=1\in\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ and $\dot{\mathbf{a}}=\mathbf{0}\in\mathcal{P}(\mathbf{0},\mathbf{a})$. Therefore, the viability condition (36) holds for $m=0$.
For $m=1$, according to (21) and (23) one has
$\partial\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}\subset\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}),$ (41)
and for $\mathbf{x}\in\partial\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}$, the projection $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ is unique and the circle $\mathcal{B}_{\lVert\mathbf{x}_{\pi}\rVert}(\mathbf{x})$ intersect with $\partial\mathcal{O}_{\mathcal{W}}$ only at the location $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}}).$ Hence, for all $\xi\in\mathcal{F}_{1}\setminus\mathcal{J}_{1}$ with $\mathbf{x}\in\partial\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}$,
$\mathbf{T}_{\mathcal{F}}(\xi)=\mathcal{P}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\times\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})\times\mathcal{P}(\mathbf{0},\mathbf{a})\times\{0\}\times\mathbf{T}_{\mathbb{R}_{\geq 0}}(s),$ (42)
where $\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$ and $\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ are defined in (37) and (38), respectively. Also, according to (11a), for $m=1$, one has $\mathbf{u}(\xi)=\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{a}),\kappa_{r}>0.$ From (14), it follows that $\eta(\mathbf{x})=1$ for $\mathbf{x}\in\partial\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}$. As a result, according to (11a) and (12), the control vector is simplified to $\mathbf{u}(\xi)=\kappa_{r}{\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}}.$ It can be shown that for any $\mathbf{a}\in\mathbb{S}^{2}$, the inner product $\mathbf{x}_{\pi}^{\intercal}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$ is always non-negative. Therefore, for $\xi\in\mathcal{F}_{1}\setminus\mathcal{J}_{1}$ with $\mathbf{x}\in\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}$, one has $\mathbf{u}(\mathbf{x},\mathbf{h},\mathbf{a},1,s)^{\intercal}(\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}}))\geq 0$. Moreover, according to (31), it is clear that $\dot{\mathbf{h}}=\mathbf{0}\in\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$, $\dot{s}=1$ and $\dot{\mathbf{a}}=\mathbf{0}\in\mathcal{P}(\mathbf{0},\mathbf{a})$. Hence, the viability condition in (36) holds for $m=1.$

### v4

@@ I INTRODUCTION
Safe autonomous robot navigation consists in steering a robot to a target location while avoiding obstacles. One common method is the Navigation Function (NF) approach [1], where a combination of attractive and repulsive vector fields, with a properly tuned parameter, guarantees safe, almost [^1] Almost global convergence refers to convergence from all initial locations except a set of zero Lebesgue measure. global convergence of the robot to the target location in sphere worlds. For general convex and star-shaped obstacles, diffeomorphic mappings [1, 2] allow the use of the NF approach but require complete knowledge of the environment. In [3], the NF method has been extended to convex obstacles that satisfy certain curvature properties, and are not too flat and not too close to the target. The work in [4] removes this flatness limitation, in environments with ellipsoidal obstacles, by using the Hessian to locally transform areas near obstacles into spherical regions. However, similar to [3], it is assumed that the entire shape of the obstacle is known when the robot visits its neighborhood.
@@
In [10], a feedback controller based on Nagumo’s theorem [11, Theorem 4.7] was proposed, guaranteeing safe, almost global convergence to the target location in a priori unknown environments with strongly convex obstacles. The work in [12] introduced a reactive navigation approach using separating hyperplanes for robots in environments with unknown, sufficiently separated, strongly convex obstacles.
However, the approaches discussed above provide, at best, almost global convergence guarantees due to the undesired equilibria that are generated when using continuous time-invariant vector fields [13]. This can be resolved by introducing discontinuities in the control, as shown in [14, 15, 16, 17].
In [14], a discontinuous feedback law was developed for robot navigation in partially known two-dimensional environments, guiding the robot with the navigation function gradient near known obstacles and boundary-following near unknown ones. In [15], the authors proposed a discontinuous feedback controller for autonomous navigation of nonholonomic robots in two-dimensional environments with non-convex obstacles, imposing restrictions on the curvatures of obstacles’ boundaries and the inter-obstacle arrangements. In [16], the authors proposed a hybrid control law to globally asymptotically stabilize a class of linear systems with drift while avoiding neighborhoods of unsafe isolated points. In [17], hybrid control techniques were employed to achieve global stabilization of target locations in $n$-dimensional environments with sufficiently separated ellipsoidal obstacles.
@@ II-A Notations
For two sets $\mathcal{A},\mathcal{B}\subset\mathbb{R}^{n}$, the relative complement of $\mathcal{B}$ with respect to $\mathcal{A}$ is denoted by $\mathcal{A}\setminus\mathcal{B}=\{\mathbf{a}\in\mathcal{A}|\mathbf{a}\notin\mathcal{B}\}$. The symbols $\partial\mathcal{A},\mathcal{A}^{\circ}$, $\mathcal{A}^{c}$ and $\bar{\mathcal{A}}$ represent the boundary, interior, complement and the closure of the set $\mathcal{A}$, respectively, where $\partial\mathcal{A}=\bar{\mathcal{A}}\backslash\mathcal{A}^{\circ}$. The cardinality of a set $\mathcal{A}$ is denoted by $\mathbf{card}(\mathcal{A})$. The Minkowski sum of the sets $\mathcal{A}$ and $\mathcal{B}$ is denoted by, $\mathcal{A}\oplus\mathcal{B}=\{\mathbf{a}+\mathbf{b}|\mathbf{a}\in\mathcal{A},\mathbf{b}\in\mathcal{B}\}$. The dilated version of a set $\mathcal{A}\subset\mathbb{R}^{n}$ with $r\geq 0$ is represented by $\mathcal{D}_{r}(\mathcal{A})=\mathcal{A}\oplus\mathcal{B}_{r}(\mathbf{0})$. The $r-$neighborhood of a set $\mathcal{A}$ is denoted by $\mathcal{N}_{r}(\mathcal{A})=\mathcal{D}_{r}(\mathcal{A})\setminus\mathcal{A}^{\circ}$ where $r$ is a strictly positive scalar.
@@ II-D Hybrid system framework
where the (set-valued) flow map $\mathbf{F}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ and jump map $\mathbf{J}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ govern continuous and discrete evolution, which can occur, respectively, in the flow set $\mathcal{F}\subset\mathbb{R}^{n}$ and the jump set $\mathcal{J}\subset\mathbb{R}^{n}$. The notions of solution $\phi$ to a hybrid system, hybrid time domain dom $\text{dom }\phi$, maximal and complete solution are, respectively, as in [20, Def. 2.6, Def. 2.3, Def. 2.7, p.30].
@@ III PROBLEM FORMULATION
We consider a spherical robot with radius $r\geq 0$ operating in a convex, three-dimensional Euclidean space $\mathcal{W}\subseteq\mathbb{R}^{3}.$ The workspace $\mathcal{W}$ is cluttered with finite number of compact, convex obstacles $\mathcal{O}_{i}\subset\mathcal{W},i\in\{1,\ldots,b\},b\in\mathbb{N}.$ We define obstacle $\mathcal{O}_{0}:=(\mathcal{W}^{\circ})^{c}$ as the complement of the interior of the workspace. Collectively, $\mathcal{O}_{\mathcal{W}}=\bigcup_{i\in\mathbb{I}}\mathcal{O}_{i}$, where $\mathbb{I}=\{0,\ldots,b\}.$ We make the following workspace feasibility assumption:
@@ IV-B Geometric construction of the flow and jump sets
When the robot operates in the move-to-target mode, its velocity is directed towards the target location. Hence, if the path joining the robot’s location and the target is obstructed by an obstacle, let’s say $\mathcal{O}_{i}$ for some $i\in\mathbb{I}$ i.e., $\mathcal{L}_{s}(\mathbf{x},\mathbf{0})\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\neq\emptyset$, then the center of the robot enters the $(r_{a}+\gamma)-$neighborhood of the obstacle $\mathcal{O}_{i}$ through the landing region $\mathcal{R}_{l}^{i},$ which is defined as
@@ IV-B1 Flow and jump sets (move-to-target mode)
where $\gamma_{s}\in(0,\gamma).$ For robustness purposes (with respect to noise), we introduce a hysteresis region by allowing the robot, operating in the move-to-target mode inside the $(r_{a}+\gamma)-$neighborhood of the set $\mathcal{O}_{\mathcal{W}}$, to move closer to the set $\mathcal{O}_{\mathcal{W}}$ before switching to the obstacle-avoidance mode.
@@ IV-B2 Flow and jump sets (obstacle-avoidance mode)
The robot operates in the obstacle-avoidance mode inside the $\gamma-$neighborhood of the obstacle-occupied workspace. Since the robot can safely move straight towards the target location and exit the $\gamma-$neighborhood of the nearest obstacle, whenever it is in the exit region (17), it should switch back to the move-to-target mode only if it is in the exit region. To that end, we make use of the hit point $\mathbf{h}$ (i.e., the location of the center of the robot when it switched from the move-to-target mode to the current obstacle-avoidance mode) to define the jump set of the obstacle-avoidance mode $\mathcal{J}_{1}^{\mathcal{W}}$ for the state $\mathbf{x}$ as follows:
@@ V STABILITY ANALYSIS
The next lemma shows that the hybrid closed-loop system (31) satisfies the hybrid basic conditions [20, Assumption 6.5], which guarantees the well-posedness of the hybrid closed-loop system.
@@ VII SIMULATION RESULTS
We compare our approach with the separating hyperplane method developed in [12] which can be implemented in a priori unknown environments using information from a range-bearing sensor mounted on the robot. However, unlike our approach, it only works for strongly convex obstacles [12, Assumption 2]. The workspace contains two obstacles, an ellipsoidal obstacle (strongly convex) and a polygonal obstacle that does not satisfy [12, Assumption 2]. The robot’s radius is set to $0.2m$, with a minimum safety distance $r_{s}=0.1m$ and $\gamma=0.5m$. The gains $\kappa_{s}=0.4$ and $\kappa_{r}=2$ are used in (11a), while the sensing radius $R_{s}=2m$ is used in (33), and $\epsilon=0.5m$ in (22).
@@ IX-B Proof of Lemma 2
IX-C Proof sketch for Lemma 3
Since the proposed hybrid closed-loop system (31) satisfies the hybrid basic conditions outlined in Lemma 2, we make use of [20, Proposition 6.10] to show the forward invariance of $\mathcal{K}$. Specifically, we verify the satisfaction of the following viability condition:
@@ IX-C Proof sketch for Lemma 3
which allows us to establish the completeness of the solution $\xi$ to the hybrid closed-loop system (31). In (35), $\mathbf{T}_{\mathcal{F}}(\xi)$ represents the tangent cone to the set $\mathcal{F}$ at $\xi$.